OpenStudy (calculusxy):

MEDAL!!! A ball is dropped from a very tall building. Assume no air resistance. What is the velocity of the ball when the distance travelled is 100m?

OpenStudy (calculusxy):

@DanJS

OpenStudy (danjs):

Did you learn the kinematic equations?

OpenStudy (danjs):

\[\Delta y = V _{0}*t + \frac{ 1 }{ 2 }*g*t^2\]

OpenStudy (danjs):

delta x is 100m, The original Velocity is Zero, and the acceleration of gravity g, is 9.8m/s^2

OpenStudy (danjs):

Find how long it has fallen t... us that in Final Velocity = a*t

OpenStudy (danjs):

Or, there is the already combined version \[v _{f}^2 = v _{i}^2+2*g*\Delta y\]

OpenStudy (danjs):

just solve for final velocity, everything else is given

OpenStudy (danjs):

\[v _{f}=\sqrt{2*g*\Delta y}\]